Length of a rectangle is three times its breadth. Perimeter of the rectangle is 40cm. Find its length and width.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,

Perimeter of rectangle = 40 cm
Length of a rectangle is three time its breadth = 3b
We know that, perimeter of rectangle = 2 × (length + breadth)

40 = 2 × (3b + b) 50 = 2 × 4b
40 = 8b
40/8 = b

b = 5 cm
Therefore, width of rectangle = 5 cm

Length of rectangle = 3b
=3 × 5

= 15 cm

Thanks